# Reducing quantum error correction overhead using soft information

### v2

@@ III Discussion
Our simulations show that soft decoding increases $\Lambda$ compared to hard decoding, making it possible to accomplish the same logical error rates as a hard decoder while using fewer physical qubits. When decoding surface code quantum memory experiments with realistic physical error rates and a high amount of soft measurement flips, soft LCD improves $\Lambda$ by $(20\pm 2)\%$ for neutral atoms and by $(11\pm 1)\%$ for superconducting qubits. We also find that shortening measurement times can increase $\Lambda$ for superconducting qubits when decoding with soft information, showing that faster-than-usual QEC circuits can be compatible with good sub-threshold scaling. Furthermore, our simulations employing soft-BM and soft-BP+OSD confirm that the advantageous scaling properties of soft decoding can be replicated in different QEC codes and decoders.
@@
To fully incorporate soft information into large-scale error-corrected devices, various engineering challenges need to be overcome. The additional information of up to $8$ bits per measurement [16] necessitates higher-bandwidth hardware for communication between the readout system and the classical decoder. This can be alleviated by compression, where measurements whose probability substantially deviates from the high-confidence regions of 00 or $1$ are allocated $N$ bits, while measurements with a high-likelihood outcome are transmitted as a single bit. A second challenge is decoding speed. In the setting of graph decoders, soft decoding is only possible by dynamically updating the edge weights in a decoding graph, but graph re-initialisation at runtime may be too slow to keep up with the throughput requirements of superconducting qubits. Decoders designed for fast graph processing such as LCD or Helios [49] may be necessary to keep the decoding time within a strict time budget. This challenge also extends to BP-based decoders, where soft information updates require fast decoder architectures such as those proposed in Refs. [45, 46].
@@
While our results map the fidelity landscape in a setting where leakage is not present, further work is needed to determine if soft information about leakage events, i.e. probabilistic leakage flags determined during the classification process, can be used to decode QEC experiments on superconducting platforms in real time. While leakage can be treated as a discrete event as done in Refs. [16, 13, 21], better logical fidelities may be reached by quantifying the leakage probability as done with the neural network decoder of Ref. [8]. As graph-based decoders are likely to be required for fast-throughput and low-latency QEC experiments, it remains an open question how to incorporate probabilistic leakage updates into a graph decoding workflow.
On neutral atom qubit platforms, the readout duration is a major bottleneck in overall system speed. As seen in our work, soft information may be used to shorten measurement times while maintaining the high logical fidelity needed for effective QEC, but this is yet to be demonstrated in a real experiment. We hope future investigations can shed light on the effectiveness of soft information decoding for improving the logical fidelity and logical clock speed of error-corrected neutral atom devices.
IV Methods
IV.1 Noise models and simulation parameters
@@
Our numerical simulations of QEC circuits leverage the fast Clifford simulator Stim [33] to generate measurement samples. For each qubit platform, we add circuit-level noise according to a physically informed model as detailed in section IV.1.2 for neutral atoms and section IV.1.1 for superconducting devices. We treat the Stim-generated measurement samples as the initial outcomes $\bar{\mu}\in\{0,1\}$ and sample soft measurements $\mu$ from the platform-specific measurement response distributions $f^{(\bar{\mu}=0)}(\mu)$ and $f^{(\bar{\mu}=1)}(\mu)$. All decoding is then performed on the posterior probabilities $P(1\mid\mu)$ for the soft decoders or the hardened outcomes $\hat{\mu}$ for the hard decoders.
IV.1.1 Superconducting qubit noise model and simulations
@@
In the time-dependence analysis of section II.3.2, we use the SI1000 model as a starting point for our circuit-level noise model and incorporate the measurement time $\tau_{\textrm{M}}$ into the noise channel probabilities as follows. We maintain the same error channel probabilities for single-qubit gates, two-qubit gates and reset operations as SI1000. Time-dependence is introduced by incorporating idling noise to qubits that are not acted upon by gates, measurements or resets, with the probability of the idling error channel determined by the specific duration of each operation. We set $\tau_{\textrm{1q}}=20~{}\text{ns}$ for single-qubit gates, $\tau_{\textrm{2q}}=40~{}\text{ns}$ for two-qubit gates and $\tau_{\textrm{R}}=40~{}\text{ns}$ for resets. Based on the duration $\tau$ of each operation, single-qubit Pauli noise with components $(P_{X},P_{Y},P_{Z})$ is added to the idling qubits according to the following probabilities, as taken from Ref. [37]:
$\begin{split}P_{X}&=\frac{1}{4}\left[1-\exp\left(-\tau/\textrm{T}_{1}\right)\right]\\ P_{Y}&=\frac{1}{4}\left[1-\exp\left(-\tau/\textrm{T}_{1}\right)\right]\\ P_{Z}&=\frac{1}{2}\left[1-\exp\left(-\tau/\textrm{T}_{2}\right)\right]-\frac{1}{4}\left[1-\exp\left(-\tau/\textrm{T}_{1}\right)\right]\\ \end{split}$ (13)
@@
IV.1.2 Neutral atom noise model and simulations
@@
×$\quad\times\left[\sum_{k=0}^{\mu}\frac{(\eta R_{0}+R_{b\to d})^{k}(R_{\textrm{bg}}t)^{k}}{k!(\eta R_{0})^{k}}-e^{-(\eta R_{0}+R_{b\to d})t}\sum_{k=0}^{n}\frac{(\eta R_{0}+R_{b\to d})^{k}(\eta R_{0}+R_{\textrm{bg}})^{k}t^{k}}{k!(\eta R_{0})^{k}}\right]$ (15a)
@@
×$\quad\times\left[e^{-R_{d\to b}t}\sum_{k=0}^{\mu}\frac{(\eta R_{0}-R_{d\to b})^{k}(R_{\textrm{bg}}t)^{k}}{k!(\eta R_{0})^{k}}-e^{-\eta R_{0}t}\sum_{k=0}^{\mu}\frac{(\eta R_{0}-R_{d\to b})^{k}(\eta R_{0}+R_{\textrm{bg}})^{k}t^{k}}{k!(\eta R_{0})^{k}}\right]$ (15b)
@@
The main parameters for this readout model include the scattering rate of an atom in the bright state $R_{0}$, the detection efficiency $\eta$, the background scattering rate $R_{\textrm{bg}}$, the scattering rate when atoms transition from a bright to a dark state $R_{b\to d}$, and the rate of scattering when atoms transition from a dark to a bright state $R_{d\to b}$. In a setting where background scattering and transition errors are low, the longer the readout pulse the smaller the overlap is between the two distributions, resulting in a lower soft measurement flip probability. The distributions for the bright and the dark states are Poissonian, with an effective mean no. of photons are $\lambda_{\textrm{bright}}=(\eta R_{0}+R_{\textrm{bg}})\tau_{\textrm{M}}$ in a bright state and $\lambda_{\textrm{dark}}=R_{\textrm{bg}}\tau_{\textrm{M}}$ in a dark state. To reflect a realistic neutral atom readout chain [24], our simulations are configured with $\eta=0.1\%$, $\tau_{\textrm{M}}=100~{}\mu\textrm{s}$, $R_{0}=10^{7}$, $R_{\textrm{bg}}=10^{3}$, $R_{b\to d}=960$ and $R_{d\to b}=2$. The qubits in neutral atoms are naturally identical and have long coherence times, therefore, it is safe to assume uniform measurement response functions across all qubits in a simulation.
IV.2 Estimating the error suppression rate
@@
The following process describes how we extract the error suppression rate $\Lambda$ from surface code simulations of variable code distances $d$. The method is applicable to experiments in both $X$- and $Z$-basis. When quoting results for a $\Lambda$-factor in section II, we take the average across $\Lambda_{X}$ and $\Lambda_{Z}$. In a quantum memory experiment taking $T$ rounds of syndrome extraction, the logical error probability is given by
$P_{\textrm{L}}=\frac{1}{2}\left[(1-(1-2\epsilon_{\textrm{L}})^{T}\right]$ (16)
where $\epsilon_{\textrm{L}}$ is the per-round logical error rate [10]. Solving the above for $\epsilon_{\textrm{L}}$ in the limit of low error rates $\epsilon_{\textrm{L}}\ll 1$ as done in Ref. [16] gives the approximate relation
$\epsilon_{\textrm{L}}\approx\frac{P_{\textrm{L}}}{T}\quad.$ (17)
@@
To extract $\Lambda$ as defined in Eq. 4, we first compute the per-round error rate $\epsilon_{\textrm{L}}(d)$ from Eq. 17 for varying code distances $d$, and take a least-squares fit of $x=(d+1)/2$ against $y=\log\left(\epsilon_{\textrm{L}}(d)\right)$. The resulting fit takes the form
$\log\left[\epsilon_{\textrm{L}}(d)\right]=-\log(\Lambda)\times d+C$ (18)
@@
where $C$ is a constant offset, and we can extract $\Lambda$ from the coefficient in front of $d$. To estimate the uncertainty in $\Lambda$, we use the Sinter library (a sub-package of Stim [33]) to evaluate hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data. This method gives us error bars on the logical error probability $P_{\textrm{L}}$, avoiding the pitfall of making overconfident estimates about the accuracy of the logical error probability at low error rates. We propagate the Sinter-derived uncertainty $\Delta P_{\textrm{L}}$ to $\epsilon_{\textrm{L}}$ via Eq. 17 and use the resulting error bars in the curve fitting process of Eq. 18 to determine the standard error in $\Lambda$.
IV.3 Soft information for LCD
@@
IV.4 Soft Belief Propagation Decoder
@@
Input: Circuit-level Tanner graph $\mathcal{T}_{\textrm{circ}}$, posterior measurement probabilities $P(1\mid\mu_{i})$ for each soft measurement $\mu_{i}$, prior probabilities ${p_{\textrm{priors}}}$
@@
V Data and code availability
The Stim circuits used to simulate the experiments in this study, as well as code used to generate soft measurement samples for superconducting and neutral atom platforms are available at https://github.com/riverlane/soft_information_models.
VI Acknowledgments

### v3

@@ IV Discussion
Our simulations show that soft decoding increases $\Lambda$ compared to hard decoding, making it possible to accomplish the same logical error rates as a hard decoder while using fewer physical qubits. When decoding surface code quantum memory experiments with realistic physical error rates and a high amount of soft measurement flips, soft LCD improves $\Lambda$ by $(20\pm 2)\%$ for neutral atoms and by $(11\pm 1)\%$ for superconducting qubits. These correspond to $33\%$ and $13\%$ reductions in the physical qubit footprints required for MegaQuop-scale computation on the two platforms respectively—demonstrating that soft information can be a powerful tool for reducing the hardware overheads of error-corrected quantum computation. In addition to the physical space reductions, we also find benefits in soft information for making QEC circuits faster. Shortening measurement times can increase $\Lambda$ for superconducting qubits when decoding with soft information, showing that faster-than-usual QEC circuits can be compatible with good sub-threshold scaling. Furthermore, our simulations employing soft-BM and soft-BP+OSD confirm that the advantageous scaling properties of soft decoding as seen in [12, 16] can be replicated in different QEC codes and decoders.
@@
V Conclusion
While our simulations show that soft information offers significant advantages in the error suppression performance of QEC protocols, key aspects of our results are yet to be verified experimentally. It has been shown for superconducting qubits that soft decoding can achieve a higher $\Lambda$-factor than traditional hard decoding in a repetition code [16], but these results include a leakage-decoding component which may obscure the effects of soft information on its own. Larger-scale experiments with real-time decoding such as seen in Ref. [1] could be designed with a fast soft information decoder, offering critical insights into the scalability of soft decoding protocols for future large-scale error-corrected devices. Such work could also quantify the impact of soft decoding on decoding time, likely to be realised with a hardware decoding platform such as those proposed in Refs. [17, 57]. Experimentalists could also leverage soft information as part of the measurement calibration workflow, confirming if it can be used to tune an error-corrected device for the highest possible logical fidelity. A potentially rich area for further optimisation is the use of high-dimensional classification protocols such as neural networks [58], which may benefit from additional rich input such as the measurement time traces of a superconducting processor. Such advanced classifiers could further improve the fidelity of the measurement posteriors, enhancing decoding accuracy.
Further experiments need not be constrained to surface codes or repetition codes thanks to the introduction of bivariate bicycle codes [23]. Experimental realisations of soft decoding on atomic qubits with qLDPC codes present a major milestone in this area, and neutral atoms may be particularly well suited to near-term demonstrations of soft decoding on BB codes thanks to their high qubit connectivity. As prior implementations of soft decoding have been solely focused on solid-state qubits, employing more efficient QEC codes on atomic qubits would confirm the wide applicability of soft decoding techniques for different qubit platforms. The slower operating speeds of atoms likely make any additional time overhead resulting from soft information processing more easily manageable, enabling such QEC demonstrations to be performed in real time. It would also be of great interest to see how an $XZZX$ surface code [59] tailored to the strongly $Z$-biased noise of neutral atoms would perform against a typical square surface code, and whether soft information could be used to further augment performance in this setting.
@@ V Conclusion
In addition to experimental realisations of the protocols studied in this paper, further work is needed to determine how soft information about leakage events, i.e. probabilistic leakage flags determined during the classification process, can be used to decode QEC experiments on superconducting platforms. While leakage can be treated as a discrete event as done in Refs. [16, 13, 17], better logical fidelities may be reached by quantifying the leakage probability as done with the neural network decoder of Ref. [8]. As graph-based decoders are likely to be required for fast-throughput and low-latency QEC experiments, it remains an open question how to incorporate probabilistic leakage updates into a graph decoding workflow.
VI Appendix
VI.1 Noise models and simulation parameters
@@
Our numerical simulations of QEC circuits leverage the fast Clifford simulator Stim [40] to generate measurement samples. For each qubit platform, we add circuit-level noise according to a physically informed model as detailed in section VI.1.2 for neutral atoms and section VI.1.1 for superconducting devices. We treat the Stim-generated measurement samples as the initial outcomes $\bar{\mu}\in\{0,1\}$ and sample soft measurements $\mu$ from the platform-specific measurement response distributions $f^{(\bar{\mu}=0)}(\mu)$ and $f^{(\bar{\mu}=1)}(\mu)$. All decoding is then performed on the posterior probabilities $P(1\mid\mu)$ for the soft decoders or the hardened outcomes $\hat{\mu}$ for the hard decoders.
VI.1.1 Superconducting qubit noise model and simulations
@@
To define our experiments in section III.1.1, we fix the two-qubit depolarising error rate $p$ and the soft measurement error probability $p_{\textrm{S}}$ to our desired values. From these we derive the gate error rates as specified by SI1000, and define the measurement response functions $f^{(\bar{\mu}=0)}(\mu)$ and $f^{(\bar{\mu}=1)}(\mu)$ with a SNR that satisfies Eq. 11.
In the time-dependence analysis of section III.1.2, we use the SI1000 model as a starting point for our circuit-level noise model and incorporate the measurement time $\tau_{\textrm{M}}$ into the noise channel probabilities as follows. We again start by fixing $p$ and maintain the same error channel probabilities for single-qubit gates, two-qubit gates and reset operations as SI1000. Time-dependence is introduced by incorporating idling noise to qubits that are not acted upon by gates, measurements or resets, with the probability of the idling error channel determined by the specific duration of each operation. We set $\tau_{\textrm{1q}}=20~\text{ns}$ for single-qubit gates, $\tau_{\textrm{2q}}=40~\text{ns}$ for two-qubit gates and $\tau_{\textrm{R}}=40~\text{ns}$ for resets. Based on the duration $\tau$ of each operation, single-qubit Pauli noise with components $(P_{X},P_{Y},P_{Z})$ is added to the idling qubits according to the following probabilities, as taken from Ref. [46]:
[…]
@@
VI.1.2 Neutral atom noise model and simulations
@@
$\quad\times\left[\sum_{k=0}^{\mu}\frac{(\eta R_{0}+R_{b\to d})^{k}(R_{\textrm{bg}}t)^{k}}{k!(\eta R_{0})^{k}}-e^{-(\eta R_{0}+R_{b\to d})t}\sum_{k=0}^{n}\frac{(\eta R_{0}+R_{b\to d})^{k}(\eta R_{0}+R_{\textrm{bg}})^{k}t^{k}}{k!(\eta R_{0})^{k}}\right]$ (15a)
@@
$\quad\times\left[e^{-R_{d\to b}t}\sum_{k=0}^{\mu}\frac{(\eta R_{0}-R_{d\to b})^{k}(R_{\textrm{bg}}t)^{k}}{k!(\eta R_{0})^{k}}-e^{-\eta R_{0}t}\sum_{k=0}^{\mu}\frac{(\eta R_{0}-R_{d\to b})^{k}(\eta R_{0}+R_{\textrm{bg}})^{k}t^{k}}{k!(\eta R_{0})^{k}}\right]$ (15b)
@@
The main parameters for this readout model include the scattering rate of an atom in the bright state $R_{0}$, the detection efficiency $\eta$, the background scattering rate $R_{\textrm{bg}}$, the scattering rate when atoms transition from a bright to a dark state $R_{b\to d}$, and the rate of scattering when atoms transition from a dark to a bright state $R_{d\to b}$. In a setting where background scattering and transition errors are low, the longer the readout pulse the smaller the overlap is between the two distributions, resulting in a lower soft measurement flip probability. The distributions for the bright and the dark states are Poissonian, with an effective mean no. of photons $\lambda_{\textrm{bright}}=(\eta R_{0}+R_{\textrm{bg}})\tau_{\textrm{M}}$ in a bright state and $\lambda_{\textrm{dark}}=R_{\textrm{bg}}\tau_{\textrm{M}}$ in a dark state. To reflect a realistic neutral atom readout chain [28], our simulations are configured with $\eta=0.1\%$, $\tau_{\textrm{M}}=100~\mu\textrm{s}$, $R_{0}=10^{7}$, $R_{\textrm{bg}}=10^{3}$, $R_{b\to d}=960$ and $R_{d\to b}=2$. The qubits in neutral atoms are naturally identical and have long coherence times, therefore, it is safe to assume uniform measurement response functions across all qubits in a simulation.
VI.2 Estimating the error suppression rate
@@
The following process describes how we extract the error suppression rate $\Lambda$ from surface code simulations of variable code distances $d$. The method is applicable to experiments in both $X$- and $Z$-basis. When quoting results for a $\Lambda$-factor in section III, we take the average across $\Lambda_{X}$ and $\Lambda_{Z}$.
In a quantum memory experiment taking $T$ rounds of syndrome extraction, the logical error probability is given by Eq. 5. By repeatedly sampling and decoding syndromes from the circuit, we obtain an estimate of $P_{L}$. Solving Eq. 5 for the per-round error rate $\epsilon_{d}$ we get
$\epsilon_{d}=\frac{1}{2}\left(1-\left(1-2P_{L}\right)^{1/T}\right)$ (16)
@@
where we substitute the measured value of $P_{L}$. To find $\Lambda$, we first compute the per-round error rate $\epsilon_{d}$ from Eq. 4 for varying code distances $d$, and take a least-squares fit of $x=(d+1)/2$ against $y=\log\left(\epsilon_{d}\right)$. The resulting fit takes the form
$\log\left(\epsilon_{d}\right)=-\log(\Lambda)\times d+\log(p_{0})$ (17)
@@
where $p_{0}$ is a constant offset, and we can extract $\Lambda$ from the coefficient in front of $d$. To estimate the uncertainty in $\Lambda$, we use the Sinter library (a sub-package of Stim [40]) to evaluate hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data. This method gives us error bars on the logical error probability $P_{\textrm{L}}$, avoiding the pitfall of making overconfident estimates about the accuracy of the logical error probability at low error rates. We propagate the Sinter-derived uncertainty $\Delta P_{\textrm{L}}$ to $\epsilon_{d}$ via Eq. 16 and use the resulting error bars in the curve fitting process of Eq. 17 to determine the standard error in $\Lambda$.
VI.3 Soft information for LCD
@@
VI.4 Soft Belief Propagation Decoder
@@
Input: Circuit-level Tanner graph $\mathcal{T}_{\textrm{circ}}$, posterior measurement probabilities $P(1\mid\mu_{i})$ for each soft measurement $\mu_{i}$, prior probabilities $p_{\textrm{priors}}$} .
@@
VII Data and code availability
The Stim circuits used to simulate the experiments in this study, as well as code used to generate soft measurement samples for superconducting and neutral atom platforms are available at https://github.com/riverlane/soft_information_models. Additional data used to support the findings of this study can be made available by the corresponding authors upon reasonable request.
VIII Acknowledgments
@@ X Competing Interests
XI Copyright
This is the Accepted Manuscript version of an article accepted for publication in Quantum Science and Technology. IOP Publishing Ltd is not responsible for any errors or omissions in this version of the manuscript or any version derived from it. This Accepted Manuscript is published under a CC BY licence. The Version of Record is available online at https://doi.org/10.1088/2058-9565/ae4e55.